OpenStudy (anonymous):

Anyone good with differential equations ?? y'' + 4y' + 4y = et initial value y'(0)=2 , y(0)=1

OpenStudy (anonymous):

oh and ps its e^t

myininaya (myininaya):

first have you solve the for the homogeneous solution?

OpenStudy (anonymous):

no i havent tried i think they somehow use laplace transforms

myininaya (myininaya):

do we have to use laplace?

OpenStudy (anonymous):

no just solve as is so find characteristic equation and then sub in cuachy euler

OpenStudy (anonymous):

no just solve as is so find characteristic equation and then sub in cuachy euler

myininaya (myininaya):

i don't know this cuachy euler business

myininaya (myininaya):

i can show you how i would solve it

OpenStudy (anonymous):

okay

myininaya (myininaya):

r^2+4r+4=0 (r+2)^2=0 r=-2 so we have \[y_h=c_1e^{-2t}+c_2te^{-2t}\] now we need to find particular solution

myininaya (myininaya):

\[y_p=Ae^{t}=> y_p^{'}=Ae^{t}=>y_p^{''}=Ae^{t}\] \[Ae^t+4Ae^t+4Ae^t=9Ae^t=e^{t}=>A=\frac{1}{9}\]

myininaya (myininaya):

\[y=y_h+y_p=c_1e^{-2t}+c_2te^{-2t}+\frac{1}{9}e^{t}\]

myininaya (myininaya):

http://tutorial.math.lamar.edu/Classes/DE/HOHomogeneousDE.aspx this is an awesome site

myininaya (myininaya):

also you can apply the initial conditions right?